OpenStudy (anonymous):

A binomial distribution has a 60% rate success.There are 18 trials 1) What is the probability that there will be at least 12 success? 2) What is the probability that there will be 12 failures ? 3) What is the expected number of successes?

OpenStudy (dumbcow):

\[P(x=X) = \left(\begin{matrix}n \\ X\end{matrix}\right)p^{X}*(1-p)^{n-X}\] where n=18 p = 0.6 For 1) X>=12, For 2) X =6 \[E(X) = n*p\]

OpenStudy (dumbcow):

for which part

OpenStudy (anonymous):

3) I did 0.60*18= 10.8 the key is 10

OpenStudy (dumbcow):

ok, im getting 0.37 for 1) did you add each one, X=12, X=13 ...X=18

OpenStudy (dumbcow):

for 3) 10.8 is correct must be a rounding thing

OpenStudy (anonymous):

\[(\frac{18!}{(18-12)!12!})*0.6^{12} * 0.4^{18-12} =0.1655178\]

OpenStudy (anonymous):

sorry 1) the key is 0.25

OpenStudy (dumbcow):

correct unfortunately you have to do 13,14,15...18 too

OpenStudy (dumbcow):

oh ok

OpenStudy (anonymous):

can you show me how ?

OpenStudy (dumbcow):

i can try..looks like i got it wrong too :)

OpenStudy (dumbcow):

what did they give for 2) 0.0145 ??

OpenStudy (anonymous):

I can't get answer 0.25,the key may wrong ?

OpenStudy (anonymous):

2) I don't have a key ,but show me how you do number 2

OpenStudy (dumbcow):

12 failures means 6 successes Find P(x=6)

OpenStudy (dumbcow):

ok i did some double checking and as long as the question is worded correctly, you are right, the key must be wrong

OpenStudy (anonymous):

ty, I hope the key is wrong